OpenStudy (anonymous):

Use Newton's method to find an approximate solution for the equation x^2+x=3 Start with x0=1 and iterate Newton's method twice to find x2.

OpenStudy (anonymous):

f' = 2x + 1 f(0) = -3 x1 = x0 - f(0) / f'(0) shoot - i'm not sure thats right - i'll have to check

OpenStudy (anonymous):

I am also not sure, if you find anything about it pls post it

OpenStudy (amistre64):

newtons method is finding the equation of the tangent line; getting its root for a new x value to repeat the iteration right?

OpenStudy (amistre64):

Xn = f'(x)(x-x0)+y0 = 0

OpenStudy (anonymous):

I think so

OpenStudy (amistre64):

(1,2) Xn = (2(1)+1)(x-1)+2

OpenStudy (amistre64):

Xn = 3x - 1; Xn = 1/3 then right?

OpenStudy (anonymous):

what do you label as Xn?

OpenStudy (amistre64):

new X value; Xn just seemed appropriate

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

I coulda went with Larry, but.... i know a cucumber named Larry lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

sorry still not clear, I am just working it out, but I dont see how you get y0=2

OpenStudy (amistre64):

x^2 +x = 3 is that a typo?

OpenStudy (amistre64):

x^2 +x +3 maybe?

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

first one

OpenStudy (amistre64):

x^2 +x = 3 ; when x=1 this is simply a false statement then 1+1 = 3???

OpenStudy (anonymous):

yep but the method will tend to the right answer

OpenStudy (amistre64):

x^2 +x -3 = 0 then

OpenStudy (anonymous):

that is the same

OpenStudy (amistre64):

then remove the 0 to find some solutions

OpenStudy (amistre64):

when x = 1; y = ? (1,-1) then right?

OpenStudy (anonymous):

-1 right

OpenStudy (amistre64):

use this point for your tangent equation that follows: y-y0 = m(x-x0) y = m(x-x0)+y0

OpenStudy (amistre64):

m = f'(x)

OpenStudy (amistre64):

y = (2(1)+1)(x-1)-1

OpenStudy (anonymous):

clear

OpenStudy (amistre64):

y = 3x-4; the root is x = 4/3 then right?

OpenStudy (anonymous):

that is why i did not see how that 2 appeared

OpenStudy (amistre64):

:) right concept, wrong thought lol

OpenStudy (anonymous):

right and I should do this again to get the result, thanks for the help!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

A mathematica solution with comments. Refer to the pdf attachment.

OpenStudy (anonymous):

andras:what/where are you studying?

OpenStudy (anonymous):

pure maths, just finished 1st year

OpenStudy (anonymous):

are you on facebook? pardon me if posting this here is a breach of ettiquet

OpenStudy (anonymous):

:) I dont mind. Yes I am on facebook

OpenStudy (anonymous):

you can add me if you wish but I dont use facebook that often. Andras Honyek

OpenStudy (anonymous):

will do...I just sent you a request...neat...most people i know are glued to their fb accounts.